molar concentration

A 15.0% (w/w) NiCl2 (129.61 g/mol) solution has a density of 1.149 g/mL. Calculate the molar concentration of NiCl2 in this solution.
Answer #1
15% (w/w)=15g NiCl2/100g solution
mL solution: (100g)(mL/1.149g)= 87.03mL= 0.08703L
mol NiCl2: (15g)(mol/129.61g)=0.12mol
(0.12mol NiCl2)(2mol Cl-/1mol NiCl2)/0.08703L= 2.66M
answered by: NeedsHelp
Answer #2
Step1 Consider 1l of solution ; Mass of 1l of solution=1000x1.149=1149 g

Step2 Mass of NiCl2 = 1149x15/100

Step3 Molarity = 1149x15/100x129.61 =1.3298